OpenStudy (anonymous):

I really have a hard time graphing inequalities.. Any help? tips? My teacher says we have to put it into the y=mx+b form but I have no idea how to do that. here's a problem I dont understand how to graph or put into that form. 6x+3y>12 or 12x+y(greater than equal to) -4

OpenStudy (anonymous):

Right not 6x+3y>12 is in standard form so when you turn in to slope intercept form: y=mx+b, 1) take 6x and minus it from both sides: 3y>12-6x 2) isolate the y by dividing 3 y>4-2x When graphing inequalities like this, pretend that this sign: > is actually = then graph the line normally. Then, you shade the side that the inequality is on and since it isn't \[\ge \] or \[\le \] you have to have the line be dotted. And when you want to know which side to shade, take a coordinate for example (0,0) plug it in your equation; y>4-2x and see if it works it would be 0>4 which doesn't work so plug in another coordinate to see if it works.